Hi, please assist.Given the demand for a product as Qd = 100 - 5P and the supply is given as Qs = -20 + 3p. You are told equilibrium is obtained at the point where Qd = Qs. The equilibrium quantity for the product is

1
Expert's answer
2021-06-03T13:03:40-0400

Qd=Qs

100-5p = -20+3p

100+20=3p+5p

120/8 = 8p/8

P=15

Therefore,

Equilibrium Quantity (Qd)

Qd = 100 - 5P

=100-5(15)

=100-75

Qd = 25